OpenStudy (anonymous):

Dr. Collins borrowed some money to buy new furniture for her office. She paid $720.00 simple interest on a 7.5-year loan at 16%. Find the principal

OpenStudy (tad1):

I = Prt P = I/rt = 720/(.16)(7.5) no calculator so you'll have to do the arithmetic